Sir/Madam I want to find the value of Sin37 Trigonomtrically! Please help me! With Regards! Anish

Sir/Madam
I want to find the value of Sin37 Trigonomtrically! Please help me!
With Regards!
Anish

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
in a right angled triangle with sides as 3,4,5 or 6,8,10 , the angles other than the right angles are 37 degre and 53 degree . 
so, 
sin 37 degree = 3/5 = cos 53 degree 
sin 53 degree = 4/5 = cos 37 degree 
the value will come around 0.60181574 , 
else , u can do it by taylor series , but understanding it will be complicated . 
so, better follow this method , 
ALL THE BEST ..